OpenStudy (saifoo.khan):

3 sin x + 5 cos x = 0

OpenStudy (slaaibak):

3sinx=-5cosx 3sinx/cosx = -5 tan x = -5/3 I don't have a calculator, but you should know how to solve from there

OpenStudy (anonymous):

Subtract the 5cos(x)=>3sin(x)=-5cos(x) Divide by -5cos(x)=> (-3/5)tan(x)=1 Multiply by -5/3=>tan(x)=-5/3=>x=arctan(-5/3)

OpenStudy (anonymous):

I've done it too, and it's ok.

OpenStudy (saifoo.khan):

k, thnks guyz/

OpenStudy (anonymous):

No problem.